# Inference-time Alignment in Continuous Space

### v2

@@ 5.6 Further Analysis
Computational Complexity. As shown in Section 5.5, we compare the time and memory efficiency of SEA with the sentence-level BoN and token-level ARGS methods. All experiments were conducted under consistent settings using LLaMA-3-8B-Base, a LLaMA-3.2-3B reward model, and 2×A100 80GB GPUs. Results show that SEA matches BoN-64’s efficiency at larger Langevin steps (s=50) and surpasses it with fewer steps (s=10), while consistently achieving better performance. This advantage stems from SEA ’s candidate-free, sentence-level optimization, which avoids ARGS’s expensive token-by-token search and BoN’s reliance on large candidate sets, leading to both higher efficiency and stronger alignment.
@@
Robustness to Reward Model Quality. We also evaluate the robustness of SEA to reward model quality in Table 9 (Appendix C.2). The results show that SEA maintains strong performance even with imperfect reward models, demonstrating its robustness and practical applicability in real-world scenarios where reward signals may be noisy or imperfect.
@@ B.4 The Details of Baselines
We compare SEA with following inference-time alignment methods: BoN [19, 3] for $N=8,32,64$. CBS [22], ARGS [21] and RS [20].We tuned the hyperparameters for each baseline on each dataset and base model, and reported the best performance results. The hyperparameters search range we used are provided in  Table 6. The detailed introduction is provided as follow:
@@
Best-of-$N$ [19, 3] Method generates $N$ responses for a single prompt as candidates, and selects the response which has the best behavior of the $N$ candidates based on the evaluation of a reward model, as the final response.
@@
Alignment as Reward Guided Search [21] samples top-$k$ tokens $V^{(k)}$ for the previous context $\mathbf{x}$ with highest likelihood from the base model at each step, and for each candidate $v\in V^{(k)}$, evaluates the reward $r([\mathbf{x},v])$ based on the reward model. ARGS method computes the score with the reward coefficient $w$ of each candidate at each search step as in Equation 4. For greedy version of ARGS, a candidate will be selected as Equation 5. For stochastic version of ARGS, token is sampled from a renormalized distribution among the Top-k candidate tokens.
@@
Rejection Sampling [20] Method samples $t$ tokens for a prompt $\mathbf{x}$ from base LM as a candidate, and evaluates the reward score $r([\mathbf{x},y_{\mathrm{candidate}}])$ based on the reward model. Under soft mode, a candidate will be selected only if Equation 6 holds. Under hard mode, a candidate will be selected only if $r([\mathbf{x},y_{\mathrm{candidate}}])>\tau_{r}(t)$. The reward threshold $\tau_{r}$ is set by Equation 7, where $r^{*}$ is the final reward score aimed to achieve and $r_{0}$ is the initial reward threshold set from the reward score of the prompt $r_{\mathbf{x}}$ and $r^{*}$, where $r_{0}=(1-\alpha)\cdot r_{\mathbf{x}}+\alpha\cdot r^{*}$.
@@ B.5 The Details of Implementation
(3) Notably, a major challenge in non-autoregressive sequence optimization is the unknown sequence length, which can result in incomplete or redundant text generation. To address this, we feed the generated sentence back into the model and use a revision prompt to guide the model in refining the output, either completing unfinished sentences or removing redundancy to enhance fluency. The prompts used are provided in Appendix G.
@@ C.2 Robustness to Reward Model Quality
C.3 Additional Models, Datasets and Baselines
We conducted additional experiments with Re-Control [24], an editing-based state-of-the-art method, and expanded evaluation to the HH-RLHF benchmark (results in Table 11). Furthermore, we evaluated SEA on LLaMA2-13B-Chat, a stronger RLHF-aligned base model (results in Table 11). Across all experiments, SEA consistently outperforms competing approaches, demonstrating SEA’s robustness across alignment benchmarks and model scales.
@@ C.3 Additional Models, Datasets and Baselines
Method AdvBench HarmRate $\downarrow$ (%) TruthfulQA TruthRate (%) TruthfulQA InfoRate (%) HH-RLHF WinRate (%)
@@
Method AdvBench HarmRate $\downarrow$ (%) TruthfulQA TruthRate (%) TruthfulQA InfoRate (%) GSM8K Acc (%)
@@
C.4 Additional Results for Reward Dynamics
In this section, we present additional results on reward dynamics, as shown in Figure 9 and Figure 9, which correspond to the results on AdvBench and MATH, respectively. It can be observed that SEA effectively improves the reward score (represented by the yellow line) while reducing the reference loss (represented by the blue line). The results demonstrate that for AdvBench, both the reward score and the reference loss converge rapidly within 30–60 iterations. For MATH, the reward score stabilizes around 80 iterations, while the reference loss exhibits greater fluctuations compared to AdvBench, possibly due to the increased difficulty in optimizing reasoning steps compared to natural language. However, the reference loss eventually decreases and stabilizes.
@@ C.4 Additional Results for Reward Dynamics
C.5 Additional Results for KL Budget
In this section, we present additional results on the evolution of KL divergence between optimized and initial responses across token positions over iterations. As shown in Figure 9, the ability of SEA to balance the KL budget over potisions is not an isolated phenomenon but a consistent characteristic.
@@ Appendix E Limitations and Broader Impacts
Limitations: Our approach shares a common limitation with other inference-time alignment methods: higher inference latency compared to training-time approaches. Nonetheless, as demonstrated in Section 5.6, SEA achieves better computational efficiency than comparable alternatives. Future work could investigate learning amortized inference strategies to approximate SEA’s optimization process, potentially reducing latency while preserving inference-time alignment performance.

### v4

@@ 5.6 Further Analysis
Computational Complexity. As shown in Section˜5.5, we compare the time and memory efficiency of SEA with the sentence-level BoN and token-level ARGS methods. All experiments were conducted under consistent settings using LLaMA-3-8B-Base, a LLaMA-3.2-3B reward model, and 2×A100 80GB GPUs. Results show that SEA matches BoN-64’s efficiency at larger Langevin steps (s=50) and surpasses it with fewer steps (s=10), while consistently achieving better performance. This advantage stems from SEA ’s candidate-free, sentence-level optimization, which avoids ARGS’s expensive token-by-token search and BoN’s reliance on large candidate sets, leading to both higher efficiency and stronger alignment.
@@
Robustness to Reward Model and Reward Hacking. We evaluate the robustness of SEA to reward model quality in Table 9 (Appendix C.2) and resistance against reward hacking in Table 10 (Appendix C.3). Results show that SEA maintains strong performance even with imperfect reward models and does not exhibit more severe reward hacking compared to BoN, despite higher performance.
Robustness to Reward Model and Reward Hacking. We evaluate the robustness of SEA to reward model quality in Table 9 (Appendix C.2) and resistance against reward hacking in Table 10 (Appendix C.3). Results show that SEA maintains strong performance even with imperfect reward models and does not exhibit more severe reward hacking compared to BoN, despite higher performance.
@@ 6 Conclusion
Acknowledgment
This work was supported by the Strategic Priority Research Program of the CAS (No. XDB0680302), and the National Natural Science Foundation of China (No.U21B2046, No.62202448).
@@ B.4 The Details of Baselines
We compare SEA with following inference-time alignment methods: BoN [19, 3] for $N=8,32,64$. CBS [22], ARGS [21] and RS [20].We tuned the hyperparameters for each baseline on each dataset and base model, and reported the best performance results. The hyperparameters search range we used are provided in  Table˜6. The detailed introduction is provided as follow:
@@
Best-of-NN [19, 3] Method generates $N$ responses for a single prompt as candidates, and selects the response which has the best behavior of the NN candidates based on the evaluation of a reward model, as the final response.
@@
Alignment as Reward Guided Search [21] samples top-$k$ tokens $V^{(k)}$ for the previous context $\mathbf{x}$ with highest likelihood from the base model at each step, and for each candidate $v\in V^{(k)}$, evaluates the reward $r([\mathbf{x},v])$ based on the reward model. ARGS method computes the score with the reward coefficient $w$ of each candidate at each search step as in Equation˜4. For greedy version of ARGS, a candidate will be selected as Equation˜5. For stochastic version of ARGS, token is sampled from a renormalized distribution among the Top-k candidate tokens.
@@
Rejection Sampling [20] Method samples $t$ tokens for a prompt $\mathbf{x}$ from base LM as a candidate, and evaluates the reward score $r([\mathbf{x},y_{\mathrm{candidate}}])$ based on the reward model. Under soft mode, a candidate will be selected only if Equation˜6 holds. Under hard mode, a candidate will be selected only if $r([\mathbf{x},y_{\mathrm{candidate}}])>\tau_{r}(t)$. The reward threshold $\tau_{r}$ is set by Equation˜7, where $r^{*}$ is the final reward score aimed to achieve and $r_{0}$ is the initial reward threshold set from the reward score of the prompt $r_{\mathbf{x}}$ and $r^{*}$, where $r_{0}=(1-\alpha)\cdot r_{\mathbf{x}}+\alpha\cdot r^{*}$.
@@ B.5 The Details of Implementation
(3) Notably, a major challenge in non-autoregressive sequence optimization is the unknown sequence length, which can result in incomplete or redundant text generation. To address this, we feed the generated sentence back into the model and use a revision prompt to guide the model in refining the output, either completing unfinished sentences or removing redundancy to enhance fluency. The prompts used are provided in Appendix˜G.
@@ C.2 Robustness to Reward Model Quality
C.3 Resistance against Reward Hacking
[…]
Table 10: Resistance of SEA against Reward Hacking
Method GPT-4o WinRate (%) Qwen-Max WinRate (%) DeepSeek-R1 WinRate (%) KL($\downarrow$)
BoN-8 70.69 66.15 73.45 1.20
BoN-32 71.62 69.11 73.05 2.50
BoN-64 73.03 69.25 75.39 3.17
BoN-128 74.62 70.96 75.77 3.86
BoN-256 74.81 71.15 75.58 4.55
SEA 90.45 91.31 93.21 3.91
C.4 Additional Models, Datasets and Baselines
We conducted additional experiments with Re-Control [24], an editing-based state-of-the-art method, and expanded evaluation to the HH-RLHF benchmark (results in Table˜12). Furthermore, we evaluated SEA on LLaMA2-13B-Chat, a stronger RLHF-aligned base model (results in Table˜12). Across all experiments, SEA consistently outperforms competing approaches, demonstrating SEA’s robustness across alignment benchmarks and model scales.
@@ C.4 Additional Models, Datasets and Baselines
Method
AdvBench
HarmRate $\downarrow$ (%)
TruthfulQA
TruthRate (%)
TruthfulQA
InfoRate (%)
$H$-RLHF
WinRate (%)
@@
Method
AdvBench
HarmRate $\downarrow$ (%)
TruthfulQA
TruthRate (%)
TruthfulQA
InfoRate (%)
GSM8K
Acc (%)
@@
C.5 Additional Results for Reward Dynamics
In this section, we present additional results on reward dynamics, as shown in Figure˜9 and Figure˜9, which correspond to the results on AdvBench and MATH, respectively. It can be observed that SEA effectively improves the reward score (represented by the yellow line) while reducing the reference loss (represented by the blue line). The results demonstrate that for AdvBench, both the reward score and the reference loss converge rapidly within 30–60 iterations. For MATH, the reward score stabilizes around 80 iterations, while the reference loss exhibits greater fluctuations compared to AdvBench, possibly due to the increased difficulty in optimizing reasoning steps compared to natural language. However, the reference loss eventually decreases and stabilizes.
@@ C.5 Additional Results for Reward Dynamics
C.6 Additional Results for KL Budget
In this section, we present additional results on the evolution of KL divergence between optimized and initial responses across token positions over iterations. As shown in Figure˜9, the ability of SEA to balance the KL budget over potisions is not an isolated phenomenon but a consistent characteristic.
@@ Appendix E Limitations and Broader Impacts
Limitations: Our approach shares a common limitation with other inference-time alignment methods: higher inference latency compared to training-time approaches. Nonetheless, as demonstrated in Section˜5.6, SEA achieves better computational efficiency than comparable alternatives. Future work could investigate learning amortized inference strategies to approximate SEA’s optimization process, potentially reducing latency while preserving inference-time alignment performance.